According to the rule of 70, if a country's real GDP per capita grows at an annual rate of 5% instead of 7%, it will take how many additional years for that country to double its level of real GDP per capita? (Show Your Work)  
Answer #1

According to the rule of 70, we can determine the number of years in which an amount doubles by dividing 70 by the growth rate.

So, when the per capita income grows by 5%, the number of years required to double the per capita income = 70/5 = 14 years.

On the other hand, when the per capita income grows by 7%, the number of years required to double the per capita income = 70/7 = 10 years.

So, when the growth rate is 5%, the number of additional years required = 14 - 10 = 4 years.
Answer #2

SOLUTION :


As per rule of 70.


Years required to double the amount = 70 / growth rate in % 


Hence,


Years required to double the GDP at 7% growth rate = 70 / 7 = 10 years


Years required to double the GDP at 5% growth rate = 70 / 5 = 14 years


So, additional years required to double the GDP at 5% growth rate instead of 7% :

= 14 - 10

= 4 years (ANSWER).

answered by: Tulsiram Garg